OpenStudy (anonymous):

5 ping-pong balls with a total surface area of 320πcm² are placed side by side in a long box. Find the length, in cm of the box. (Answer given in the book is 40cm.) Please help thanks! TT

zepdrix (zepdrix):

So the length of the tube will be 5 times the diameter of the balls.

zepdrix (zepdrix):

Or in other words, 10 times the radius.

zepdrix (zepdrix):

The Surface Area equation involves the radius of the spherical object. I'll bet we can solve for the radius and then find the length! :)

zepdrix (zepdrix):

Here is our equation for surface area. \[\large A_s=4\pi r^2\] They told that the TOTAL surface area of the balls is \(320\pi\). If we divide that value by 5, we'll get the surface area for just 1 ball.

zepdrix (zepdrix):

So we can set it up like this. Here is the surface area of `one ball`.\[\large \frac{320\pi}{5}=4\pi r^2\]From here, we can solve for r to find the radius of `one ball`.

zepdrix (zepdrix):

Dividing each side by 4pi gives us,\[\large \frac{320\pi}{20\pi}=r^2\]Which simplifies to,\[\large 16=r^2\]Taking the square root gives us the radius of each ball. Then we would want to multiply by 2 to get the diameter of each ball. Then multiply by 5 to get the length of 5 balls (the length of the box).

OpenStudy (anonymous):

Thanks so much I understand now! XD